OpenStudy (anonymous):

Hello. I am having problems solving 6(3z-1)-5(z+3)=8(z+1)....please help

OpenStudy (anonymous):

first you do all your multiplication. \[18z-6-5z+15=8z+8\] this is what it would look like after the multiplication. then you would combine like terms \[13z+9=8z+8\] this is what it would look like after combining like terms then you would move the 8z to the left side of the equation and the 9 to the right of the equation \[5z=-1\] this is what it would look like after that then you would divide both ides by 5 \[z=-1/5\] this is what that would look like so your answer is z= -1/5

OpenStudy (anonymous):

the answer was 29/5....can we try a similar problem. 6(4z-1)-5(z+3)=8(z+1)